OpenStudy (anonymous):

Use an inequality to solve the problem. Be sure to show the inequality and all of your work. In order for a chemical reaction to remain stable, its Celsius temperature must be no more than 111.4 degrees Celsius. Find the Fahrenheit temperatures at which the reaction will remain stable.Use F=9/5(C)+32 So will this be correct? We were given the Celsius of the temperature needed which is not more than 111.4. We must find the Fahrenheit of the temperature so that it remains stable. Since the Celsius is less then or equal to the Fahrenheit Using F= (9/5)C + 32 , we now have F = (9/5)-111.4 +3

OpenStudy (anonymous):

F = (9/5)-111.4 +32 F = 9/5 *5/9 -111.4 + 32 F= -79.4 Check: -79.4= 9/5 -111.4 +32 -79.4=-79.4 True We found that F= -79.4 which is less than 111.4, so the temperature at which the reaction will stay stable is at -79.4 degrees Fahrenheit. The answer is -79.4 which is less then the Celsius will stabilize the reaction of the temperature.